OpenStudy (anonymous):

What is the sum of a 30–term arithmetic sequence where the first term is 73 and the last term is –101?

OpenStudy (anonymous):

a = 73 a +29d = -101 or l = -101 s = 30/2{ 73 - 101) s = -994

OpenStudy (anonymous):

the choices are -476, -448, -420, -392

OpenStudy (anonymous):

okay then i must have done some mistake

OpenStudy (anonymous):

yeah your correct sorry -420